. A toy truck of mass 10kg travels south at 2ms. It collides with a toy car of mass 9kg. Determine the initial velocity of the toy car if both objects come to rest.


1
Expert's answer
2021-02-28T07:34:48-0500

Let's write the law of conservation of momentum:


"m_1v_1+m_2v_2=0,""m_1v_1=-m_2v_2,""v_2=-\dfrac{m_1v_1}{m_2}=-\dfrac{10\ kg\cdot2\ \dfrac{m}{s}}{9\ kg}=-2.22\ \dfrac{m}{s}."

The sign minus means that the initial velocity of the toy car directed in opposite direction to the motion of the toy truck (to the North).